OpenStudy (anonymous):

In a sample of seven cars, each car was tested for nitrogen-oxide emissions (in grams per mile) and the following results were obtained: 0.08, 0.18, 0.12, 0.16, 0.14, 0.15, 0.11. Assuming that this sample is representative of the cars in use, construct a 98% confidence interval estimate of the mean amount of nitrogen-oxide emissions for all cars. If the EPA requires that nitrogen-oxide emissions be less than 0.165 g/mi, can we safely conclude that this requirement is being met? What is the confidence interval estimate of the mean amount of nitrogen-oxide emissions for all cars? __g/mi < µ < __g/mi Can we safely conclude that the requirement that nitrogen-oxide emissions be less than 0.165 g/mi is being met? a. No, it is possible that the requirement is being met, but it is also very possible that the mean is not less than 0.165 g/mi b. No, because the confidence interval does not contain 0.165 g/mi c. Yes, we can definitely conclude that the requirement is met for all cars. d. Yes, because the confidence interval contains 0.165 g/mi

OpenStudy (anonymous):

bar x = 0.08+0.12+0.16+0.14+0.15+0.11/7 = 0.94/7 = .13428 is how far I got!! Im stuck here!! Can someone please show me how to input this into the Ti84??